SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (Oct. 4, 2005)

Attendees: Rainer Bartoldus, Kris Chaisanguanthum, Al Eisner, Yury Kolomensky, Selina Li, Minghui Lu, Stephen Sekula, Dan Walker.

Operations

  • Last Tuesday (Sept. 27) owl shift the TSF-Y crate started to have trouble (see HN and the ROM logfile). This probelm was discovered not from fast mon plots but from the fact that fastmon crashed. It happened soon after DIRC had the problem with the power supply, therefore we didn't get to take much physics data on last Tuesday. Runs 58384-58421 potentially got affected by this problem although currently we only see the last 9 minutes of 58384 and all events in 58385 are bad. On Wednesday (Setp.28) swing shift, a DCC board for the TSF-Y crate was replaced by a temporary "spare" taken out from the old TSF-Y crate with help from Matt Weaver. The bad DCC board was confirmed to be bad at the test stand as well and it is being shipped to LLR in France for Antoine Mathieu to repair. Rainer suggested us to find out why fastmon crashed since it shouldn't crash production.
  • http://babar-hn.slac.stanford.edu:5090/HyperNews/get/DetOper/4460/2.html

EMT

  • It looks like an offline build 18.3.3 is going to include the EMT prescale configuration and soon be implemented for online. We might be able to install the prescale before the shutdown.
  • Al asked if there is anything else we want to prescale and if the prescale needs to be at the same event. Rainer doesn't think it's necessary to prescale at the same event. The next possible thing to prescale is ZPD, but it doesn't show up as a problem yet, so it might not be necessary.

DCT/DCZ

  • Marc Kelly had a new version of the TSF firmware for us to test at the test stand. Su Dong and Selina tested the power cycling, and it seems to be working reliably in recovering after master crate was power cycled. However, when Jamie ran a MC test with the new firmware at the test stand, he found a low level problem for the x firmware that was reproducible. Marc is working on a fix for this.

L3

  • Al suggested us to review the to-do-list for L3 during the TRG operation meeting next week (since last week we decided to have the Friday L3 meeting join the Tuesday operation meeting).

Data Quality and Monitoring

  • Kris processed runs 58385-9 by hand for the fastmon archive since they were missing due to the fastmon crashed mentioned in the operation section above.
  • We are currently validating MC test runs made with 18.6.0.

Trigger Configurations and Rates

  • See presentation from Minghui about the 1E trigger studies done for single photon physics.
  • It was suggested to look at the 1G rate for events that pass 1M. After we get an estimate of this 1G rate, they can perhaps propose a test run with triggering on 1G instead of 1E.

This page is created by Selina Li (selina@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.